# InfNeRF: Towards Infinite Scale NeRF Rendering with O(log n) Space Complexity

### v1

@@ 5 Conclusion and Future work
In this work, we introduce InfNeRF for large-scale scene reconstruction. InfNeRF employs an LoD tree to divide the scene in space and scale into cubes, which are reconstructed using many NeRFs. During rendering, InfNeRF selectively retrieves a minimal subset of nodes, significantly reducing memory requirements and I/O time for parameter retrieval from disk or cloud storage. The memory complexity for rendering with InfNeRF is notably efficient at . Another outstanding feature of InfNeRF is that parent nodes in the LoD tree act as a smoother version of the scene, effectively reducing aliasing artifacts. In our experiments, InfNeRF achieves superior rendering quality, with an improvement of over 2.4dB in PSNR while accessing only 17% of the total parameters. Furthermore, InfNeRF imposes no constraints on the underlying NeRF, allowing seamless integration with faster and smaller NeRF models in the future to enhance their scalability and anti-aliasing quality.
@@
InfNeRF demonstrates its potential for large-scale scene reconstruction. However, there are several aspects for future exploration and improvement. Firstly, the reconstruction time and computational burden still exceed traditional, well-optimized photogrammetry methods [30, 14], necessitating further performance enhancements. Secondly, exploring the combination of octree from diverse image sources, such as satellites, planes, and drones, at different times and scales, could enable the creation of larger scenes. With these works, digitizing the Earth using NeRF will become a compelling possibility for the future.

### v2

@@ 5. Conclusion and Future work
In this work, we introduce InfNeRF for large-scale scene reconstruction. InfNeRF employs an LoD tree to divide the scene in space and scale into cubes, which are reconstructed using many NeRFs. During rendering, InfNeRF selectively retrieves a minimal subset of nodes, significantly reducing memory requirements and I/O time for parameter retrieval from disk or cloud storage and also reducing aliasing artifacts. In our experiments, InfNeRF achieves superior rendering quality, with an improvement of over 2.4 dB in PSNR while accessing only 17% of the total parameters.
@@
InfNeRF demonstrates its potential for large-scale scene reconstruction. However, there are several aspects for future exploration and improvement. Firstly, the reconstruction time and computational burden still exceed traditional, well-optimized photogrammetry methodsĀ (Schonberger and Frahm, 2016; DJI, 2024), necessitating further performance enhancements. Secondly, exploring the combination of octree from diverse image sources, such as satellites, planes, and drones, at different times and scales, could enable the creation of larger scenes. With these works, digitizing the Earth using NeRF will become a compelling possibility for the future.